OpenStudy (anonymous):

Please Help! on's new mini computer monitor has a 20 cm by 24 cm LCD screen with a plastic trim border. The total area of the monitor and border is 672 cm2. How wide is the trim around the monitor?

OpenStudy (anonymous):

(20+w)(24+w)=672 480+44w+w^2=672 w^2 + 44w - 192 = 0 (w+48)(w-4) = 0 w = -48 or w = 4 w > 0 so w = 4 the trim is half the extra part wide, so the trim is 2 cm wide!

OpenStudy (anonymous):

dang your the best!